Circular No. 6291

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

GRO J1744−28

S. Sazonov and R. Sunyaev, Space Research Institute, Moscow, on behalf of the Granat/WATCH Team, report: “On Jan. 14–15 the Galactic Center region was observed by the Granat/WATCH instrument. At the edge of the field of view, a strong flux, 4.4 ± 0.3 Crab (8–20 keV), was detected from the direction of the new transient x-ray burster and pulsar GRO J1744−28, discovered by BATSE. The WATCH data show that the source’s spectrum is much flatter in the standard x-ray band than in the BATSE sensitivity range. It is similar to typical spectra of accreting x-ray pulsars. At the distance of the Galactic Center, the luminosity of GRO J1744−28 in the 8–20 keV band is ∼ 4 × 10^{38} erg/s, a value that exceeds the Eddington luminosity for a neutron star. However, super-Eddington luminosity was observed earlier in SMC X-1 and several other x-ray pulsars. GRANAT/WATCH will continue to observe GRO J1744−28 in scanning mode, while observational conditions are improving daily.”

J. Swank, NASA Goddard Space Flight Center, reports, for the X-Ray Timing Explorer instrument and operations team: “XTE observed the new galactic transient GRO J1744−28 (IAUC 6272, 6275, 6276, 6284, 6285, 6286, 6290) for 11 hours on Jan. 18–19. A source pulsing at 2.1 Hz, as well as nine bursts 2–8 times brighter than the persistent emission, was observed from a position within the error box reported by Hurley et al. (IAUC 6286). The XTE best-fit position, from scans across the source, is α = 17^{h}44^{m}.5±0^{m}.2, δ = −28^{o}45′ ±2′ (equinox 2000.0), where the errors are systematic. A sinusoidal pulse shape was observed, and the burst flux was sometimes strongly modulated. Following each burst, the flux is depressed and takes a few minutes to recover. The spectra of both the pulsed and the burst flux are characteristic of bright accreting pulsars; they are consistent with a power law with number index ∼ 1.2 and a rollover above 13 keV with an e-folding energy of 14 keV. The absorbing column density is < 10^{22} cm^{−2}. The low column density implies that the source is probably closer than the Galactic Center and that at least the non-burst luminosity probably does not exceed the Eddington limit. The non-burst flux was ∼ 2×10^{−7} erg cm^{−2} s^{−1} (2–100 keV). These observations were taken during the period of initial instrument verification, so the calibrations are preliminary. See

http://heasarc.gsfc.nasa.gov/docs/xte/SOF/TOONews.html for fur-

ther information about results, data and future observations.”

1996 January 21  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1996 CBAT  Brian G. Marsden